Stokes Theorem - Vector

Verify Stokes theorem for F =(y^2 + x^2 - x^2)i + (z^2 + x^2 - y^2)j + (x^2 + y^2 - z^2)k over the portion of the surface x^2 + y^2 -2ax + az = 0
Answer #1

Given lean is x²ty²-29x+ az zo The equation of the given surface can be written as (x-aj? + y2 = -a (Z-a) This shows that thethe same boundary c both s . also and curl s have F = 1 रिप- Ty?t?? x²+x2-y? x²+² 7² L = 2cy-z î + 207-21 ſ + 2(x-ysk abo ñ